# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ 5.2 Convergence Rate of Algorithm 2
$=\mathcal{O}(\frac{\gamma\kappa_{g}}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu).$ (71)
where the last equality follows from $1-\rho_{D}=\Omega(\mu)$, $1-\beta=\Omega(1/\kappa_{g})$, and $L_{y\mu}=\mathcal{O}(\kappa_{g})$. Now combining the primal and dual gap functions we conclude that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\gamma}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu+\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\kappa_{g}^{3}+\kappa_{g}^{2}/\mu}{\tau K}})$. Selecting $\mu=\mathcal{O}(\kappa_{g}^{3/5}/K^{1/5})$ and $\gamma=\mathcal{O}(\kappa_{g}^{9/5}/K^{3/5})$, one can verify that the primal step-size $\tau$ can be selected as a constant since $\tau\leq\mathcal{O}(\mu^{2}/(\gamma\kappa_{g}^{3}))$ which implies that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\kappa_{g}^{3/5}}{K^{1/5}}+\frac{\kappa_{g}^{16/5}}{K^{2/5}}+\frac{\kappa_{g}^{3}}{\sqrt{K}})$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 2 requires at least $K=\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}}+\frac{\kappa_{g}^{8}}{\epsilon^{2.5}}+\frac{\kappa_{g}^{6}}{\epsilon^{2}})$ iterations.
@@
Suppose the premises of Theorem 5.4 holds and we further assume that $L_{yy}^{\Phi}=0$. Let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{2})$ then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{1}{\epsilon^{4}}+\frac{\kappa_{g}^{4}}{\epsilon^{2}})$ iterations.
6 Numerical Experiments
In this section, we evaluate the performance of our algorithms in solving the SPB problem. Specifically, we implement our algorithms to address a robust multi-task linear regression example as described in Section 1.3. We utilize both real-world and synthetic datasets, comparing the convergence results of the proposed algorithms across these datasets. For all algorithms, the step-sizes are selected based on the suggestions provided by the theoretical evaluations and are adjusted for enhanced performance across each dataset considered {NLSY79, MTL-dataset, synthetic dataset}. In particular, For i-BRPD:OPF, we let $\gamma=\{0.3,0.7,0.3\}\times\frac{1}{K^{2/3}}$, and $\mu=\{0.1,0.1,0.05\}\times\frac{1}{K^{1/3}}$; for i-BRPD:FP, we let $\gamma=\{1,9,0.05\}\times\frac{1}{K^{1/2}}$, $\mu=\{0.01,0.2,0.01\}\times\frac{1}{K^{1/4}}$, and $\tau=\{0.7,0.7,0.7\}$; in both proposed algorithms, we consider $\alpha=\eta=\frac{2}{\mu_{g}+L_{g}}$, and $\sigma=\frac{2}{2\mu}$; for MORBiT [20], we let $\sigma=\{1,0.01,0.001\}\times\frac{1}{K^{3/5}}$, $\alpha=\{1,1,1\}\times\frac{1}{K^{2/5}}$, $\tau=\{0.1,0.1,0.01\}\times\frac{1}{K^{3/5}}$. In all algorithms, we set the maximum number of iterations $K=10^{4}$.
6.1 Robust Multi-task Linear Regression
@@
In this section, we investigate the performance of our proposed algorithms for solving a robust multi-task learning Problem (1.3). To do this, we consider a robust multi-task linear regression example. In particular, each task $i\in\{1,\ldots,N\}$ consists of a linear regression with training and validation datasets denoted by $\mathcal{D}_{i}^{tr}$ and $\mathcal{D}_{i}^{val}$, respectively. For each task $i$, we consider a linear regression model in the form of $b_{i}=A_{i}(\lambda_{i}y_{i}+(1-\lambda_{i})x)+\epsilon_{i}$, which is a linear combination of the common coefficients $(x)$ and task-specific coefficients $(y_{i})$ using a given dataset $\mathcal{D}_{i}=(A_{i},b_{i})$, where $A_{i}\in\mathbb{R}^{n_{i}\times d}$ and $b_{i}\in\mathbb{R}^{n_{i}}$. We define the loss function as $\ell_{i}(y_{i};\mathcal{D}_{i})=\frac{1}{2n_{i}}\left\|A_{i}y_{i}-b_{i}\right\|^{2}$ and assume the feature space is shared among all tasks. This problem can be formulated as follows:
$\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{N}}\max_{i=1}^{N}~{}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})\equiv\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{N}}\max_{\eta\in\Delta_{N}}\sum_{i=1}^{N}\eta_{i}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})$
@@
where we encounter a strongly convex lower-level (LL) problem, while the upper-level (UL) problem is a nonconvex-concave saddle point (SP) problem similar to our main Problem (1).
To showcase the performance of the proposed algorithms in solving Problem (6.1) across various datasets, we examine three regression-based datasets. Firstly, we employ the 1979 National Longitudinal Survey of Youth (NLSY79) dataset [^1] NLSY79 dataset overview ($n=6213,d=21$). Next, we utilize the MTL-dataset ($n=5642,d=14$), which comprises a mixture of datasets including bodyfat, housing, mg, mpg, and space-ga from the LIBSVM library [7]. Additionally, we include a synthetic dataset ($n=5000,d=100$), generated randomly from a standard Normal distribution. For these datasets, we conduct experiments with $N=5$ tasks. We distribute the data points evenly among the considered tasks. Then, we allocate 75% of the data for each task as the training dataset and reserve the remaining 25% as the validation dataset. To evaluate the performance of our proposed algorithms (i-BRPD:OPF and i-BRPD:FP), we compared them with the Multi-Objective Robust Bilevel Two-timescale optimization (MORBiT) algorithm introduced by [20] using three distinct datasets. As depicted in Figure 1, our proposed algorithms exhibit a faster convergence rate compared to the MORBiT algorithm, aligning with our theoretical findings. Specifically, the i-BRPD:OPF algorithm, which is a one-sided projection-free method, demonstrates the fastest convergence behavior among all algorithms. This can be attributed to the utilization of the Linear Minimization Oracle (LMO) for updating the primal variable of the objective function, circumventing the need for complex projections at each iteration as required by projection-based methods. Moreover, our fully projected gradient-based approach, i-BRPD:FP, surpasses the projection-based MORBiT algorithm. This superiority stems from our proposed algorithm’s utilization of an inexact approach for estimating the Hessian inverse matrix at each iteration, as discussed in Section 3. In contrast, the MORBiT method approximates the Hessian inverse using the Neumann series, as discussed in [15], which is less accurate and computationally expensive compared to our approach.
@@ 7 Conclusions
In this paper, we introduced a class of constrained saddle point problems with a bilevel structure (SPB), where the upper-level objective function is nonconvex-concave and smooth over a compact and convex constraint set, subject to a strongly convex lower-level objective function. By leveraging regularization and nested approximation techniques, we initially devised a bilevel primal-dual one-sided projection-free algorithm (i-BRPD:OPF), requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary point. Subsequently, we developed a bilevel primal-dual fully projected algorithm (i-BRPD:FP), capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. Our numerical experiments revealed the superior performance of our proposed algorithms compared to existing methods, which predominantly rely on projection-based approaches to address nonconvex strongly-concave upper-level objective functions or upper-level objective functions that are linear in the dual variable.

### v3

@@ 5.2 Convergence Rate of Algorithm 2
$=\mathcal{O}(\frac{\gamma\kappa_{g}}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu),$
where the last equality follows from $1-\rho_{D}=\Omega(\mu)$, $1-\beta=\Omega(1/\kappa_{g})$, and $L_{y\mu}=\mathcal{O}(\kappa_{g})$. Now combining the primal and dual gap functions, we conclude that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\gamma\kappa_{g}}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu+\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\gamma\tau\kappa_{g}^{3}}{\mu}})$. Selecting $\mu=\mathcal{O}(\kappa_{g}^{3/5}/K^{1/5})$, $\gamma=\frac{\mu^{3}}{\tau\kappa_{g}^{3}}=\mathcal{O}(1/(\tau\kappa_{g}^{6/5}K^{3/5}))$, and the primal step-size $\tau$ selected as a constant $\tau\leq\mathcal{O}(1/\kappa_{g}^{3})$ implies that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\kappa_{g}^{3/5}}{K^{1/5}})$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 2 requires at least $K=\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}})$ iterations.
@@
Suppose the premises of Theorem 5.4 holds and we further assume that $\Phi$ is strongly concave in $y$. Let $\gamma=\mathcal{O}(\epsilon^{2})$ and $\tau=\mathcal{O}(1/\kappa_{g}^{3})$, then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{4}})$ iterations.
Proof.
Suppose that $\Phi(x,\cdot)$ is strongly concave for any $x$ with modulus $\bar{\mu}>0$. This means that we no longer need to add a regularization parameter $\mu$ simplifying the bound on the primal and dual gap functions obtained in Theorem 5.4. In particular, one can replace $\mu$ with $\bar{\mu}$ in the bounds of Lemmas 4.5, 4.6, and 4.7. Therefore, the bound on the primal gap function in Theorem 5.4 can be rewritten as $\mathcal{G}_{\mathcal{X}}(x_{k^{*}},y_{k^{*}})=\frac{1}{\tau}\left\|s_{k^{*}}-x_{k^{*}}\right\|\leq\sqrt{\frac{2(f(x_{0})-f(x^{*}))}{\gamma\tau(K-\bar{K})}}+\sqrt{\frac{2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+4\tau M\Gamma_{4}}{\tau(K-\bar{K})}}+\sqrt{8\tau\gamma M^{2}B_{4}}$. Next, we restate the dual gap function in (5.2). In fact for any $k\geq 1$,
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2}{\sigma}\rho_{D}^{k}\left\|y_{0}-y^{*}_{\bar{\mu}}(x_{0})\right\|+\frac{2\gamma L_{y\bar{\mu}}}{\sigma\bar{\mu}(1-\rho_{D})}D_{\mathcal{X}}+2L_{y\theta}^{\Phi}\beta\left(\frac{\rho_{D}^{k/2}}{1-\beta}+\frac{\beta^{k/2}}{1-\rho_{D}}\right)\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\frac{2\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\rho_{D})(1-\beta)}D_{\mathcal{X}}+\frac{L_{y\bar{\mu}}}{\bar{\mu}\sigma}\gamma D_{\mathcal{X}}+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\frac{\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{1-\beta}D_{\mathcal{X}}$
$=\mathcal{O}\left(\max\{\rho_{D},\beta\}^{k/2}+\gamma\kappa_{g}^{2}\right).$
Therefore, $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(\gamma\kappa_{g}^{2}+\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\gamma\tau\kappa_{g}^{3}})}$, choosing $\gamma=\mathcal{O}(\kappa_{g}^{1.5}/K^{1/2})$ and the primal-step-size as constant $\tau=\mathcal{O}(1/\kappa_{g}^{3})$, it follows that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(\frac{\kappa_{g}^{3/4}}{K^{1/4}}+\frac{\kappa_{g}^{3.5}}{K^{1/2}})}$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 2 requires $K={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{4}})}$ iterations.
Remark 5.3.
We would like to highlight that our results in Theorem 5.4 demonstrate a new bound on the primal and dual gap function defined in Definition 2.2. As shown in Corollary 5.5, Algorithm 2 achieves an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations, presenting a novel result under the general setting for nonbilinear SPB problems. This matches the best-known result of $\mathcal{O}(\epsilon^{-5})$ from [20, 21], which is achieved under the specific problem setup where the upper-level objective function is linear in the maximization component. Furthermore, when the upper-level function $\Phi$ is strongly concave in $y$, our result improves to $\mathcal{O}(\epsilon^{-4})$ which is the best-known result for the deterministic setting. In the stochastic setting, [24] achieves an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations for the implicit gradient method, where they used a stronger notion of stationarity compared to ours. To achieve the same $\epsilon$-stationary point, they incur an additional complexity factor of $\mathcal{O}(\epsilon^{-2})$, resulting in an overall complexity of $\mathcal{O}(\epsilon^{-6})$ in the stochastic setting – see Proposition 4.11 in [33].
Remark 5.4.
According to the results of Corollaries 5.2 and 5.5, we have demonstrated that for solving the general non-bilinear SPB problem, our proposed methods i-BRPD:OPF and i-BRPD:FP require $\mathcal{O}(\epsilon^{-4})$ and $\mathcal{O}(\epsilon^{-5})$ iterations to achieve an $\epsilon$-stationary point, respectively. However, one needs to compare these bounds with extra caution as the metrics used to measure the gap function and the oracle used by the algorithm (i.e., LMO and PO) are fundamentally different and are not directly comparable.
6 Numerical Experiment
@@
In this section, we conduct two experiments to evaluate the performance of our algorithms in solving the SPB problem by comparing them against the state-of-the-art method MORBiT [21]. Specifically, we implement our algorithms to address a robust multi-task linear regression example as described in Section 1.3. In particular, each task $i\in\{1,\ldots,T\}$ consists of a linear regression model with training and validation datasets denoted by $\mathcal{D}_{i}^{tr}$ and $\mathcal{D}_{i}^{val}$, respectively. For each task $i$, we consider a model in the form of $b_{i}=A_{i}(\lambda_{i}y_{i}+(1-\lambda_{i})x)+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,s)$, which includes a common coefficient $x\in\mathbb{R}^{d}$, task-specific coefficients $y_{i}\in\mathbb{R}^{d}$, and parameter $\lambda_{i}\in[0,1]$ for a given dataset $\mathcal{D}_{i}=(A_{i},b_{i})$ where $A_{i}\in\mathbb{R}^{n_{i}\times d}$ and $b_{i}\in\mathbb{R}^{n_{i}}$. We define the loss function as $\ell_{i}(y_{i};\mathcal{D}_{i})=\frac{1}{2n_{i}}\left\|A_{i}y_{i}-b_{i}\right\|^{2}$.
In our experiments, we consider three datasets: the 1979 National Longitudinal Survey of Youth (NLSY79) dataset [^1] NLSY79 dataset overview ($n=6213,d=21$), the MTL-dataset ($n=5642,d=14$), which comprises a mixture of datasets including bodyfat, housing, mg, mpg, and space-ga from the LIBSVM library [8], and a synthetic dataset ($n=5000,d=100$), generated randomly from a standard Gaussian distribution, where $n$ represents the number of data points, and $d$ denotes the number of features in each dataset. We distribute the data points evenly among $N=5$ tasks. Then, we allocate 75% of the data for each task as the training dataset and reserve the remaining 25% as the validation dataset. To evaluate the performance of our proposed algorithms, we plot the iterates in terms of the gap function defined in Definition 2.1.
For all algorithms, we let the maximum number of iterations $K=10^{4}$, and the step-sizes are selected based on their theoretical results. In particular, for i-BRPD:OPF, we let $\gamma=\nu/K^{2/3}$, $\mu=\nu/K^{1/3}$; for i-BRPD:FP, we let $\gamma=\nu/K^{1/2}$, $\mu=\nu/K^{1/4}$, and $\tau=0.7$; in the both proposed algorithms, we consider $\alpha=\eta=\frac{2}{\mu_{g}+L_{g}}$, and $\sigma=\frac{1}{\mu}$; for MORBiT, we let $\sigma=\nu/K^{3/5}$, $\alpha=1/K^{2/5}$, and $\tau=\nu/K^{3/5}$, where $\nu>0$ is a tuning parameter that is adjusted to observe the best performance for the considered methods.
Experiment 1.
Considering the worst-task training, we aim to solve the following SPB problem:
$\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{T}}\max_{\eta\in\Delta_{T}}\sum_{i=1}^{T}\eta_{i}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})$ (31)
@@
where the constraint $\|x\|_{1}\leq Q$ is imposed to induce sparsity for some $Q>0$. Note that due to the separability of the lower-level problems in $y_{i}$, (31) can be formulated as a special case of (1) by considering $(x,\lambda)\in\mathbb{R}^{d+T}$ as the minimization variables, $\eta\in\mathbb{R}^{T}$ as the maximization variable, $\theta=[y_{i}]_{i=1}^{T}\in\mathbb{R}^{dT}$ as the lower-level variable. In this case, $g((x,\lambda),\theta)=\sum_{i=1}^{T}\ell_{i}(\lambda_{i}y_{i}+(1-\lambda_{i})x;\mathcal{D}_{i}^{tr})+\frac{\rho}{2}\left\|\theta\right\|^{2}$, and $\Phi((x,\lambda),\theta,\eta)=\sum_{i=1}^{T}\eta_{i}\ell_{i}(y_{i};\mathcal{D}_{i}^{val})$. It is worth mentioning that, to compare our algorithms with the MORBiT method, we consider a special case of (1) where the upper-level objective function is assumed to be linear in the maximization component.
The performance of the algorithms is depicted in Figure 1. Notably, our proposed algorithms exhibit a faster convergence compared to MORBiT. This superiority stems from using an inexact nested approximation technique for estimating the Hessian inverse matrix in our proposed method (see the discussion in Section 3). In contrast, MORBiT approximates the Hessian inverse using the Neumann series which entails a biased approximation requiring multiple Hessian evaluations. Moreover, i-BRPD:OPF algorithm, which is a one-sided projection-free method, demonstrates the fastest convergence behavior among all algorithms, possibly due to allowing for a larger selection of step-size parameters.
Experiment 2.
We conduct an additional experiment to explore a more general setting in which $\Phi$ is nonlinear in the maximization variable. Specifically, we examine the robust multi-task linear regression problem, similar to the previous experiment but with a different uncertainty set. In particular, consider the following problem:
$\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{T}}\max_{\eta\in\mathcal{P}}\sum_{i=1}^{T}\eta_{i}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})$
[…]
where $\mathcal{P}\triangleq\{\eta\in\Delta_{T}:V(\eta,\frac{1}{T}\mathbb{1}_{T})\leq\rho^{\prime}\}$ and $V(\cdot,\cdot)$ denotes the divergence measure between the unknown probability distribution $\eta$ and the uniform distribution $\frac{1}{T}\mathbb{1}_{T}$ over the dataset. Using a penalty parameter $\beta>0$, we can relax the divergence constraint and obtain the following problem:
$\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{T}}\max_{\eta\in\Delta_{T}}\sum_{i=1}^{T}\eta_{i}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})-\frac{\beta}{T}\Big{(}\frac{1}{2}\|T\eta-\mathbb{1}_{T}\|^{2}-\rho^{\prime}\Big{)}$ (32)
[…]
Note that in (32) the upper-level function is nonlinear in $\eta$.
As shown in Figure 2, our proposed algorithms demonstrate fast convergence when the upper-level objective function is nonlinear in both the primal and dual variables. This underscores the effectiveness of our methods in handling general settings, demonstrating greater versatility compared to existing approaches, which may be restricted to specific problem setups.
@@
(a) NLSY97 dataset
(b) MTL dataset
(c) Synthetic dataset
Figure 2: Comparing the performance of our proposed algorithms i-BRPD:OPF (blue) and i-BRPD:FP (red) for solving robust multi-task linear regression problems when the upper-level objective function is nonlinear
@@ 7 Conclusions
In this paper, we introduced a class of constrained saddle point problems with a bilevel structure, where the upper-level objective function is nonconvex-concave and smooth over a compact and convex constraint set, subject to a strongly convex lower-level objective function. By leveraging regularization and nested approximation techniques, we initially devised a primal-dual one-sided projection-free algorithm (i-BRPD:OPF), requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary point, moreover, when the objective function in the upper-level is linear in the maximization component, our results improve to $\mathcal{O}(\epsilon^{-3})$. Subsequently, we developed a bilevel primal-dual fully projected algorithm (i-BRPD:FP), capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. This result improves to $\mathcal{O}(\epsilon^{-4})$ when the upper-level objective function is strongly concave in $y$. Our proposed methods and their convergence guarantee present new results appearing for the first time in the literature for the considered setting, to the best of our knowledge.